OpenStudy (anonymous):

Completely factor the expression 112x² – 567y²

OpenStudy (anonymous):

first of all you can notice that 112=7*16 and 567=7*81

OpenStudy (anonymous):

7(16x^2-81y^2) now 16x^2=(4x)^2 81y^2=(9y)^2

OpenStudy (anonymous):

Now you only need to know that (a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

7(4x-9y) (4x+9y)